This question is part of the post-lab question of the Recrystallization of Aspirin experiment.

This is for organic chemistry. Please be detailed, thank you!

Assuming that 1g of aspirin dissolves in 450mL of water at 10, how much aspirin would be lost in the 1.5mL of water added to the reaction mixture (in the experiment you did in the lab) if the mixture were at 10 during filtration?
Answer #1

Since 450 mL of water dissolves 1 g of aspirin, we use proportionality to find the amount that dissolves in 1.5 mL.

This gives us: 450 / 1 = 1.5 / x

Solving, we get: 1.5 / 450 = 3.33 * 10-3 g = 3.33 mg